There are currently issues with accessing /fs/ess/ locations using ondemand.

Known issues

Unresolved known issues

Known issue with an Unresolved Resolution state is an active problem under investigation; a temporary workaround may be available.

Resolved known issues

A known issue with a Resolved (workaround) Resolution state is an ongoing problem; a permanent workaround is available which may include using different software or hardware.

A known issue with Resolved Resolution state has been corrected.

Known Issues

Title Category Resolution Description Posted Updated
issue with OnDemand 6:09 - 8:39 pm Resolved

OnDemand, epi accounting queries, the Viper DB, the Medline DB, the Eweld DB,... Read more

8 years 17 hours ago 8 years 17 hours ago
my.osc.edu logins failing Account Management Resolved

Logins to my.osc.edu are failing. This is unrelated to our InfiniBand issue; a router change at OARnet is the believed cause. They are working on re-establishing the necessary routing.

8 years 1 week ago 8 years 1 week ago
Emergency InfiniBand Shutdown (All systems) Network Resolved

We have returned to service. It appears that we have resolved the networking issues enough to allow jobs to run safely. We will continue working with our vendors to fix any remaining hardware... Read more

8 years 1 week ago 8 years 1 week ago
Lustre is still offline. HPC systems back up Maintenance Resolved

Day One of the scheduled downtime has been completed, and HPC operations have resumed. As planned, Lustre work will extend into Day Two. Jobs using /fs/lustre or $PFSDIR cannot run until this work... Read more

8 years 1 month ago 8 years 1 month ago
Certain modules not accessible Software Resolved

Certain modules are not working for all clusters since the downtime.  We have reports specifically that Amber, Gaussian, and Turbomole are not working.  We are working to resolve the issue, but... Read more

8 years 1 month ago 8 years 1 month ago
Login Shell Issues on Oakley Account/Shell Resolved

UPDATE: The shells have all been switched back for affected users, and you can submit jobs normally again.  Additionally, if you are still logged in and have the incorrect shell, logging back out... Read more

8 years 1 month ago 8 years 1 month ago
Account changes temporarily suspended Account Management Resolved

We are still experiencing some account problems related to Thursday's issue. As a result, we have taken my.osc.edu offline and cannot process email changes or password resets, either via self-... Read more

8 years 1 month ago 8 years 1 month ago
Oakley Login Node Issues Login Problems Resolved

Currently users connecting via SSH to Oakley may recieve "connection refused" or "connection failed" errors if they were not logged in before this occurred.  Glenn is currently functioning... Read more

8 years 1 month ago 8 years 1 month ago
Cannot login to clusters Resolved

As of around 3PM today (Thursday 6/12), we have reports of users being unable to login in to the clusters.  The error message given will make it sound like your password is incorrect, although it... Read more

8 years 2 months ago 8 years 2 months ago
ARMSTRONG is offline Outage Resolved

ARMSTRONG is experiencing an unexpected outage. We are working on a resolution.

8 years 4 months ago 8 years 4 months ago

Pages